### v1

# Towards Domain Adaptive Neural Contextual Bandits

Abstract
Contextual bandit algorithms are essential for solving real-world decision making problems. In practice, collecting a contextual bandit’s feedback from different domains may involve different costs. For example, measuring drug reaction from mice (as a source domain) and humans (as a target domain). Unfortunately, adapting a contextual bandit algorithm from a source domain to a target domain with distribution shift still remains a major challenge and largely unexplored. In this paper, we introduce the first general domain adaptation method for contextual bandits. Our approach learns a bandit model for the target domain by collecting feedback from the source domain. Our theoretical analysis shows that our algorithm maintains a sub-linear regret bound even adapting across domains. Empirical results show that our approach outperforms the state-of-the-art contextual bandit algorithms on real-world datasets.
1 Introduction
Contextual bandit (CB) algorithms have shown great promise for naturally handling exploration/exploitation trade-off problems with optimal statistical properties. Notably, LinUCB [1] and its various adaptations [2, 3, 4, 5, 6, 7, 8, 9], have been shown to be able learn the optimal strategy when all data come from the same domain. However, these methods fall short when applied to data from a new domain. For example, a drug reaction prediction model trained by collecting feedback from mice (the source domain) may not work for humans (the target domain).
So, how does one effectively explore a high-cost target domain by only collecting feedback from a low-cost source domain, e.g., exploring drug reaction in humans by collecting feedback from mice or exploring real-world environments by collecting feedback from simulated environments? The challenge of effective cross-domain exploration is multifaceted: (1) The need for effective exploration in both the source and target domains depends on the quality of representations learned in the source domain, which in turn requires effective exploration; this leads to a chicken-and-egg dilemma. (2) Aligning source-domain representations with target-domain representations is nontrivial in bandit settings, where ground-truth may still be unknown if the action is incorrect. (3) Balancing these aspects – effective exploration and accurate alignment in bandit settings – is also nontrivial.
To address these challenges, as the first step, we allow our method to simultaneously perform effective exploration and representation alignment, leveraging unlabeled data from both the source and target domains. Interestingly, our theoretical analysis reveals that naively doing so leads to sub-optimal accuracy/regret (verified by our empirical results) and naturally leads to additional terms in the target-domain regret bound. We then follow the regret bound derived from our analysis to develop an algorithm that adaptively collect feedback from the source domain while aligning representations from the source and target domains. Our contributions are outlined as follows:
  • We identify the problem of contextual bandits across domains and propose domain-adaptive contextual bandits (DABand) as a the first general method to explore a high-cost target domain while only collecting feedback from a low-cost source domain.
  • Our theoretical analysis shows that our method can achieve a sub-linear regret bound in the target domain.
  • Our empirical results on real-world datasets show our DABand significantly improve performance over the state-of-the-art contextual bandit methods when adapting across domains.
2 Related Work
Contextual Bandits. In the realm of adaptive decision-making, contextual bandit algorithms, epitomized by LinUCB [1], have carved a niche in efficiently balancing the exploitation-exploration paradigm across a spectrum of use cases, notably in complex adaptive systems such as recommender systems [1, 10]. These algorithmic frameworks, along with their myriad adaptations [2, 3, 4, 7, 5, 6, 9, 8, 11], have decisively outperformed conventional bandit models that operate devoid of contextual awareness [12]. This superiority is underpinned by theoretical models, paralleling the insights in [12], where LinUCB variants are validated to conform to optimal regret boundaries in targeted scenarios [13]. However, a discernible gap in these methodologies is their reduced efficacy in a new domain, particularly when getting feedback involves high cost. Our proposed DABand bridges this gap by adeptly aligning both source-domain and target-domain representations in the latent space, thereby reducing performance drop when transfer across different domains.
Domain Adaptation. The landscape of domain adaptation has been extensively explored, as evidenced by a breadth of research [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. The primary objective of these studies has been the alignment of source and target domain distributions to facilitate the effective generalization of models trained on labeled source data to unlabeled target data. This alignment is conventionally attained either through the direct matching of distributional statistics [15, 26, 27, 20, 24] or via the integration of an adversarial loss [28, 29, 30, 19, 31, 21, 22, 23]. The latter, known as adversarial domain adaptation, has surged in prominence, bolstered by its theoretical foundation [32, 33, 19, 34], the adaptability of end-to-end training in neural network architectures, and its empirical efficacy. However, these methods only work in offline settings, and assumes complete observability of labels in the source domain. Therefore they are not applicable to our online bandit settings.
Domain Adaptation Related to Bandits. There is also work related to both domain adaptation and bandits. Specifically, Guo et al. [35] propose a domain adaptation method using a bandit algorithm to select which domain to use during training.
We note that their goal and setting is different from our DABand’s. (1) [35] focuses on improving accuracy in a typical, offline domain adaptation setting. In contrast, our DABand focuses on minimizing regret in an online bandit setting. (2) [35] assumes complete access to ground-truth labels in the source domain, while in our bandit setting, ground-truth may still be unknown if the action is incorrect. Therefore their work is not applicable to our setting.
3 Theory
In this section, we formalize the problem of contextual bandits across domains and derive the corresponding regret bound. We then develop our DABand inspired by this bound in Sec. 4. All proofs of lemmas, theorems can be found in the Appendix.
[…]
3.1 Preliminaries
Typical Contextual Bandit Setting. Suppose we have $N$ samples from an unknown domain $\mathcal{D}$, which is represented as $\{{\bf x}_{i,a}^{\mathcal{D}},a_{i}^{*}\}_{i\in[N],a\in[K]}$. Here, ${\bf x}_{i,a}\in\mathcal{D}$ denotes the context for action $a\in\mathcal{A}_{i}$, and $a_{i}^{*}\in\mathcal{A}_{i}$ is the ground-truth action that will receive the optimal reward. We denote as $r_{i,a}$ the received reward in round $i$ after performing action $a$ and $\widehat{a}_{i}\in\mathcal{A}_{i}$ the chose action by a bandit algorithm in round $i$. The goal in typical contextual bandits is to learn a policy of choosing actions $\widehat{a}_{i}$ in each round to minimize the regret after $N$ rounds:
$R=\sum\nolimits_{i=1}^{N}r_{i,a_{i}^{*}}-\sum\nolimits_{i=1}^{N}r_{i,\widehat{a}_{i}}.$ (1)
LinUCB. LinUCB [1] is a classic method for the typical contextual bandit setting. It assumes that the reward is linear to its input context, i.e., ${\bf x}_{i,a}$. During each round $i$, the agent selects an arm according to historical contexts and their corresponding rewards, $r_{i,a}$:
$\widehat{a}_{i}=\mathop{\rm argmax}_{a\in\mathcal{A}_{i}}\left\{{\bf x}_{i,a}^{T}\theta_{i}+\alpha\left\|{\bf x}_{i,a}\right\|_{A_{i-1}^{-1}}\right\},$
where $A_{i-1}=\gamma{\bf I}_{d}+\sum\nolimits_{j=1}^{i-1}{\bf x}_{j,\widehat{a}_{j}}{\bf x}_{j,\widehat{a}_{j}}^{T}$ with $\gamma>0$ and ${\bf x}_{j,\widehat{a}_{j}}^{T}$ as the historical selected action’s context. ${\bf x}_{i,a}$ is the context for the candidate action $a$ in round $i$. $\alpha>0$ is a hyperparameter that adjusts the exploration rate.
LinUCB with Representation Learning. In this paper, we use Neural LinUCB [11, 10] as a backbone model to handle high-dimensional context with representation learning. We assume that there exist a ground-truth encoder $\phi^{*}$ to encode the raw context ${\bf x}_{i,a}$ into a latent-space representation (encoding) $\phi^{*}({\bf x}_{i,a})$. Subsequently, the reward for the context ${\bf x}_{i,a}$ is then $r({\bf x}_{i,a})=\langle\theta^{*},\phi^{*}({\bf x}_{i,a})\rangle$ plus some stochastic noise $\epsilon$. Furthermore, we use the same setting as in [11] where the ground-truth rewards are restricted to the range of $[0,1]$. Additionally, we further bound predicted rewards by normalizing $\|\widehat{\theta}\|=1$ and $\|\widehat{\phi}({\bf x}_{i,a})\|=1$.
In the single-domain typical setting, the goal is to learn an encoder $\widehat{\phi}$ and the contextual bandit parameter $\widehat{\theta}$, such that our estimated reward $\widehat{{r}}({\bf x}_{i,a})=\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a})\rangle$ can be close to the ground-truth reward $r({\bf x}_{i,a})$. One can then use this to form derive policies to minimize the regret in Eqn. 1.
For simplicity, in Definition 3.1 below, we further denote as $f_{\mathcal{D}}$ the (ground-truth) labeling function for domain $\mathcal{D}$ and as $h\in\mathcal{H}$ a hypothesis such that $f$ is parameterized by $\phi_{\mathcal{D}}^{*},\theta_{\mathcal{D}}^{*}$, and $h$ is parameterized by $\widehat{\phi},\widehat{\theta}$.
3.2 Our Cross-Domain Contextual Bandit Setting
Typical contextual bandits operate in a single domain ${\mathcal{D}}$. In contrast, our cross-domain bandit setting involves a source domain $\mathcal{S}$ and a target domain $\mathcal{T}$. In this setting, one can only collect feedback (reward) from the source domain, but not from the target domain. Specifically: (1) We assume a low-cost source domain (experiments on mice), where for each round $i$, one has access to the contexts $\{{\bf x}_{i,a}^{\mathcal{S}}\}_{a\in[K]}$ for each candidate actions, chooses one action $\widehat{a}_{i}$, and receives reward $r^{{\mathcal{S}}}_{i,\widehat{a}_{i}}$. (2) Additionally, we assume a high-cost target domain where collecting feedback (reward) is expensive (experiments on humans); therefore, one only has access to the contexts $\{{\bf x}_{i,a}^{\mathcal{T}}\}_{a\in[K]}$ for each candidate actions for each round, but cannot collect feedback (reward) $r^{{\mathcal{T}}}_{i,\widehat{a}_{i}}$ for any action in the target domain. The goal in our cross-domain contextual bandits is to learn a policy of choosing actions $\widehat{a}_{i}$ in the target domain to minimize the target-domain zero-shot regret for $N$ (hypothetical) future rounds:
$R_{{\mathcal{T}}}=\sum\nolimits_{i=1}^{N}r^{{\mathcal{T}}}_{i,a_{i}^{*}}-\sum\nolimits_{i=1}^{N}r^{{\mathcal{T}}}_{i,\widehat{a}_{i}}.$ (2)
Difference between Eqn. 1 and Eqn. 2. In the typical setting, Eqn. 1 uses different updated policies for each round; this is also true for the source domain. In contrast, the target regret in Eqn. 2 use the same fixed policy obtained from the source domain for all $N$ (hypothetical) rounds because one cannot collect feedback from the target domain. See the difference between Definition 3.4 and Definition 3.5 below for a more formal comparison.
3.3 Formal Definitions of Error
Definition 3.1 (Labeling Function and Hypothesis).
We define the labeling function $f_{\mathcal{D}}$ and the hypothesis $h$ for domain ${\mathcal{D}}$ as follows:
$f_{\mathcal{D}}({\bf x}_{i,a}^{\mathcal{D}})$ $={r}({\bf x}_{i,a}^{\mathcal{D}})=\langle\theta_{\mathcal{D}}^{*},\phi_{\mathcal{D}}^{*}({\bf x}_{i,a}^{\mathcal{D}})\rangle+\epsilon_{i}$
$h({\bf x}_{i,a}^{\mathcal{D}})$ $=\widehat{{r}}({\bf x}_{i,a}^{\mathcal{D}})=\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a}^{\mathcal{D}})\rangle+\epsilon_{i}$
where $\theta_{\mathcal{D}}^{*}$ is the optimal predictor for domain ${\mathcal{D}}$, $\widehat{\theta}$ denotes the estimated predictor, and $\epsilon_{i}$ is the random noise.
Next, we analyze how well our hypothesis $h$ estimates the labeling function $f$ (i.e., ground-truth hypothesis). Definition 3.2 below quantifies the closeness between two hypotheses by calculating the absolute difference in their estimated rewards.
Definition 3.2 (Estimated Error between Two Hypotheses).
Assuming all contexts ${\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}}$ are from domain $\mathcal{D}$, the error between two hypotheses $h_{1},h_{2}\in\mathcal{H}$ on domain ${\mathcal{D}}$ given selected actions $\{\widehat{a}_{i}\}_{i\in[N]}\in[K]$ is
$\epsilon_{\mathcal{D}}(h_{1},h_{2})$ $=\sum\nolimits_{i=1}^{N}\Bigg{(}\big{|}h_{1}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})-h_{2}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\big{|}\Bigg{)}=\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta}_{1},\widehat{\phi}_{1}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\rangle-\langle\widehat{\theta}_{2},\widehat{\phi}_{2}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{D}})\rangle\Big{|}\Bigg{)}.$ (3)
Note that the domain ${\mathcal{D}}$ above can be the source domain ${\mathcal{S}}$ or the target domain ${\mathcal{T}}$. We then define the error of our estimated hypothesis in these domains.
Definition 3.3 (Source- and Target-Domain Error).
With $N$ samples from the source domain $\mathcal{S}$, and $f_{S}$ denoting the labeling function for ${\mathcal{S}}$, the source-domain error is then
$\epsilon_{\mathcal{S}}(f_{\mathcal{S}},h)$ $=\sum\nolimits_{i=1}^{N}\Bigg{(}\big{|}f_{S}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})-h({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\big{|}\Bigg{)}=\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)},$
where ${\bf x}_{i,\cdot}^{\mathcal{S}}$ comes from $\mathcal{S}$. Furthermore, $\widehat{a}_{i}=\mbox{arg}\max_{a}h({\bf x}_{i,a}^{\mathcal{S}})$ and $a_{i}^{*}=\mbox{arg}\max_{a}f_{S}({\bf x}_{i,a}^{\mathcal{S}})$. For simplicity, we shorten the notation $\epsilon_{\mathcal{S}}(f_{\mathcal{S}},h)$ to $\epsilon_{\mathcal{S}}(h)$ and similarly use $\epsilon_{\mathcal{T}}(h)$ for domain $\mathcal{T}$. Assuming ${\bf x}_{i,\cdot}^{\mathcal{T}}$ comes from $\mathcal{\mathcal{T}}$, the estimated error for the target domain is then:
$\epsilon_{\mathcal{T}}(f_{\mathcal{T}},h)$ $=\sum\nolimits_{i=1}^{N}\Bigg{(}\big{|}f_{\mathcal{T}}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})-h({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\big{|}\Bigg{)}=\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\Bigg{)}.$
3.4 Source Regret and Target Regret
Below define the regret for the source and target domains.
Definition 3.4 (Source Regret).
Assuming ${\bf x}_{i,\cdot}^{\mathcal{S}}$ comes from domain $\mathcal{S}$, the source regret (i.e., the regret in the source domain) is
$\mathit{R}_{\mathcal{S}}$ $=\sum\nolimits_{i=1}^{N}\Bigg{(}f_{\mathcal{S}}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})-f_{\mathcal{S}}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\Bigg{)}=\sum\nolimits_{i=1}^{N}\Bigg{(}\langle\theta_{S}^{*},\phi_{\mathcal{S}}^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta_{\mathcal{S}}^{*},\phi_{S}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Bigg{)}.$ (4)
The goal in our cross-domain contextual bandits is to learn a policy of choosing actions $\widehat{a}_{i}$ in the high-cost target domain (e.g., human experiments) by collecting feedback only in the source domain (e.g., mouse experiments). Formally, we would like to minimize the target regret as defined below.
Definition 3.5 (Target Regret and Problem Formulation).
Denoting the estimated hypothesis as $\widehat{h}=\left\{\widehat{\phi},\widehat{\theta}\right\}$, the target regret we aim to minimize is defined as
$\mathit{R}_{\mathcal{T}}$ $=\sum\nolimits_{i=1}^{N}\Bigg{(}f_{\mathcal{T}}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})-f_{\mathcal{T}}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\Bigg{)}$
[…]
where $A^{\mathcal{S}}=\gamma{\bf I}+\sum\nolimits_{i=1}^{N}[\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})][\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})]^{T}$ denotes the context matrix accumulated by the selected context features in the source domain.
3.5 Cross-Domain Error Bound for Regression
Prior to introducing our final regret bound, it is necessary to define an additional component. A fundamental challenge in general domain adaptation problems is to manage the divergence between source and target domains. Unfortunately, previous domain adaptation theory only covers classification tasks [36, 19]. In contrast, the problem of contextual bandits is essentially a reward regression problem. To address this challenge, we introduce the following new definition to bound the error between the source and target domains.
Definition 3.6 ($\mathcal{H}\Delta\mathcal{H}$ Hypothesis Space for Regression).
For a hypothesis space $\mathcal{H}$, the symmetric difference hypothesis space $\mathcal{H}\Delta\mathcal{H}$ is the set of hypotheses s.t.
$g\in\mathcal{H}\Delta\mathcal{H}\quad\iff\quad g(x)=|h({\bf x})-h^{\prime}({\bf x})|.$
We then can further define the Divergence for $\mathcal{H}\Delta\mathcal{H}$ Hypothesis Space:
$\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$ $=2\enspace\sup_{h,h^{\prime}\in\mathcal{H}}|\mathbb{E}_{x\sim\mathcal{S}}[|h(x)-h^{\prime}(x)|]-\mathbb{E}_{x\sim\mathcal{T}}[|h(x)-h^{\prime}(x)|]|$
Definition 3.7 (Optimal Hypothesis).
The optimal hypothesis, denoted as $h^{*}$, can balance the estimated error for both source and target domain. Formally,
$h^{*}$ $=\mbox{arg}\min_{h\in\mathcal{H}}\enspace\epsilon_{\mathcal{S}}(h)+\epsilon_{\mathcal{T}}(h),$
and we define the minimum estimated error for the optimal hypothesis $h^{*}$ as
$\psi$ $=\epsilon_{\mathcal{S}}(h^{*})+\epsilon_{\mathcal{T}}(h^{*}).$
Algorithm 1 DABand Training Algorithm
1:  Input: regularization parameter $\gamma>0$, number of total steps $N$, episode length $H$, exploration parameters $\alpha$.
2:  Output: parameters of the model: $\widehat{\phi}_{N},\widehat{\theta}_{N},A_{N},b_{N}$.
3:  Initialization: $A_{0}=\gamma{\bf I},b_{0}={\bf 0}$, and for $\widehat{\theta}_{0}$ and $\widehat{\phi}_{0}$ are initialized following [11].
4:  for $i=1,...,N$ do
5:     Obtain $\left\{{\bf x}_{i,a}^{\mathcal{S}}\right\}_{a\in[K]}$ and $\left\{{\bf x}_{i,a}^{\mathcal{T}}\right\}_{a\in[K]}$ from the source and the target domain, respectively.
6:     Choose an action $\widehat{a}_{i}=\mathop{\rm argmax}_{a\in[K]}\left[\widehat{\phi}_{i-1}({\bf x}_{i,a}^{\mathcal{S}})\right]\widehat{\theta}_{i-1}+\alpha\|\left[\widehat{\phi}_{i-1}({\bf x}_{i,a}^{\mathcal{S}})\right]\|_{A_{i-1}}$.
7:     Get the reward $r_{i}=\mathbf{r}({\bf x}_{i,\widehat{a}_{i}})$ based on the selected action $\widehat{a}_{i}$.
8:     Update bandit parameters:
9:     $\begin{cases}A_{i}=A_{i-1}+\left[\widehat{\phi}_{i-1}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\right]\left[\widehat{\phi}_{i-1}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\right]^{T},\\ b_{i}=b_{i-1}+r_{i}\left[\widehat{\phi}_{i-1}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\right],\\ \widehat{\theta}_{i}=A_{i}^{-1}b_{i}\end{cases}$
10:     if $mod(i,H)=0$ then
11:        for $j=1,...,H$ do
12:           Calculate $\mathcal{L}^{DABand}$ in Eqn. 10.
13:           Update encoder $\widehat{\phi}_{i}$ and discriminator $g$ by back-propagation to solve the minimax optimization in Eqn. 9.
14:        end for
15:     else
16:        $\widehat{\phi}_{i}=\widehat{\phi}_{i-1}$
17:     end if
18:  end for
19:  Output: $\widehat{\phi}_{N},\widehat{\theta}_{N},A_{N},b_{N}$.
3.6 Final Regret Bound
With all the definitions of source/target regret in Definition 3.4 and Definition 3.5, we can then bound the target regret using Theorem 3.1 below.
Theorem 3.1 (Target Regret Bound).
Denoting the contexts from the source domain ${\mathcal{S}}$ as $\{{\bf x}_{t,a}^{\mathcal{S}}\}_{i\in[N],a\in[K]}$, the associated ground-truth action as $\{a_{i}^{*}\}_{i=1}^{N}$, and the contexts from the target domain ${\mathcal{T}}$ as $\{{\bf x}_{t,a}^{\mathcal{T}}\}_{i\in[N],a\in[K]}$, the upper bound for our target regret $R_{{\mathcal{T}}}$ is
$\mathit{R}_{\mathcal{T}}$ $\triangleq\sum_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\Bigg{)}$
$\leq{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}}_{Predicted~{}Reward}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}},$ (5)
where $\mathit{R}_{\mathcal{S}}$, $\epsilon_{\mathcal{S}}(h)$, and $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$ are defined in previous sections. $\psi$ and $C$ are constants which are independent of the problems.
Here we provide several observations (for more analysis, please refer to the Appendix):
  • (1)
    Since $R_{\mathcal{S}}$ is sub-linear w.r.t. the number of samples in the source domain, so is $R_{\mathcal{T}}$.
  • (2)
    Theorem 3.1 bounds the target regret using the source regret, enabling the exploration of the target domain by only collecting feedback (reward) from the source domain.
  • (3)
    Typical generalization bounds in DA naively minimize the source regret $R_{{\mathcal{S}}}$ while aligning source and target data in the latent space (i.e., minimizing $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$). This is not sufficient, as we need two additional terms, i.e., the regression error $\epsilon_{\mathcal{S}}(h)$ and the predicted reward $\sum\nolimits_{i=1}^{N}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}$, as regularization terms.
  • (4)
    Minimizing the regression error $\epsilon_{\mathcal{S}}(h)$ encourages accurate prediction of source rewards.
  • (5)
    Minimizing the predicted reward $\sum\nolimits_{i=1}^{N}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}$ regularizes the model to avoid overestimating rewards.
Note that Theorem 3.1 is nontrivial. While it does resemble the generalization bound in domain adaptation, there are key differences. As mentioned in Observation (3) above, our target regret bound includes two additional crucial terms not found in domain adaptation. Specifically:
  • Regression Error in the Source Domain. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$ (i.e., $\epsilon_{\mathcal{S}}(h)$ in Theorem 3.1), which defines the difference between the true reward from selecting action $\widehat{a}_{i}$ and the estimated reward for this action.
  • Predicted Reward. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$, which serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Details for more analysis are in Appendix F.3.1.
4 Method
With Theorem 3.1, we can then design our DABand algorithm to obtain optimal target regret. We discuss how to translate each term in Eqn. 5 to a differential loss term in Sec. 4.1 and then put them together in Sec. 4.2.
4.1 From Theory to Practice: Translating the Bound in Eqn. 5 to Differentiable Loss Terms
Source Regret. Inspired by Neural-LinUCB [11], we use LinUCB [1] to update $\widehat{\theta}$ and the following empirical loss function when updating the encoder parameters $\widehat{\phi}$ by back-propagation in round $i$:
$\mathcal{L}_{i}^{\mathit{R}_{\mathcal{S}}}=\sum\nolimits_{a=1}^{K}\left(\widehat{\theta}^{T}\widehat{\phi}({\bf x}_{i,a})-{r}({\bf x}_{i,a})\right)^{2},$ (6)
where $\theta$ is the contextual bandit parameter, and $\widehat{\phi}$ is the encoder shared by the source and target domains, i.e., we set $\widehat{\phi}_{{\mathcal{S}}}=\widehat{\phi}_{{\mathcal{T}}}$ during training.
Insights. From Eqn. 6, a fundamental difference between classification problems and bandit problems becomes apparent. In classification tasks, both the ground-truth label and the estimated label are accessible. However, in bandit problems, we only know the estimated label and reward. If the estimated rewards align with our estimated label’s correctness, it suggests knowledge of the ground-truth label for that feature. If not, we can only deduce that the estimated label is inaccurate, without identifying the correct label among the other candidates. This uncertainty significantly amplifies the complexity of contextual bandit problems.
Regression Error. We use $L_{1}$ loss to optimize the regression error, defined as the difference between the true reward from selecting action $\widehat{a}_{i}$ and the estimated reward for this action (see Definition 3.3 for a detailed explanation of this regression error). Minimizing the regression error term directly encourages accurate prediction of source rewards using $L_{1}$ loss. Specifically, we use
$\mathcal{L}_{i}^{\epsilon_{\mathcal{S}}(h)}=\left|\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|,$ (7)
where $\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle$ is the reward received by the agent in the source domain. Minimizing the regression error term above directly encourages accurate prediction of source rewards.
Data Divergence. The data divergence term in Eqn. 5 leads to following loss term:
$\mathcal{L}^{div}=\max_{g}N\cdot(\mathbb{E}^{\mathcal{S}}[\mathcal{L}_{D}(g(\widehat{\phi}({\bf x})),0)]+\mathbb{E}^{\mathcal{T}}[\mathcal{L}_{D}(g(\widehat{\phi}({\bf x})),1)]),$
where $\mathbb{E}^{\mathcal{S}}$ and $\mathbb{E}^{\mathcal{T}}$ denote expectations over the data distributions of $({\bf x},a)$ in the source and target domains, respectively. $g$ is a discriminator that classifies whether ${\bf x}$ is from the source domain or the target domain. $\mathcal{L}_{D}$ is the binary classification accuracy, where labels 00 and $1$ indicate the source and target domains, respectively. In practice, we use the cross-entropy loss as a differentiable surrogate loss. As in [28], solving the minimax optimization $\min\nolimits_{\widehat{\phi}}\max\nolimits_{g}{\mathcal{L}}^{div}$ is equivalent to aligning source- and target-domain data distributions in the latent (encoding) space induced by the encoder $\widehat{\phi}$.
Table 1: Accuracy on the target domain for the DIGIT dataset. Note that in this zero-shot target regret setting, the accuracy $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is the target regret.
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
Test Accuracy 0 1 2 3 4 5 6 7 8 9 Average
LinUCB [1] 0.3667$\pm$0.03 0.4045$\pm$0.02 0.5102$\pm$0.03 0.3811$\pm$0.01 0.3157$\pm$0.04 0.3445$\pm$0.01 0.4229$\pm$0.03 0.4595$\pm$0.03 0.3489$\pm$0.02 0.2449$\pm$0.01 0.3808$\pm$0.02
LinUCB-P [1] 0.3645$\pm$0.02 0.6398$\pm$0.04 0.3151$\pm$0.02 0.2709$\pm$0.03 0.2258$\pm$0.01 0.1995$\pm$0.02 0.1752$\pm$0.01 0.3786$\pm$0.04 0.1273$\pm$0.02 0.2992$\pm$0.03 0.3060$\pm$0.02
NLinUCB [11] 0.3781$\pm$0.04 0.3228$\pm$0.03 0.3569$\pm$0.03 0.3381$\pm$0.04 0.3663$\pm$0.04 0.4312$\pm$0.05 0.4766$\pm$0.05 0.4004$\pm$0.05 0.4034$\pm$0.02 0.3613$\pm$0.03 0.3816$\pm$0.04
NLinUCB-P [11] 0.8588$\pm$0.03 0.2224$\pm$0.02 0.3880$\pm$0.03 0.3029$\pm$0.02 0.2978$\pm$0.03 0.0000$\pm$0.01 0.1869$\pm$0.02 0.2243$\pm$0.01 0.2148$\pm$0.01 0.0000$\pm$0.00 0.2706$\pm$0.02
DABand (Ours) 0.6879$\pm$0.10 0.6654$\pm$0.03 0.6056$\pm$0.01 0.5694$\pm$0.05 0.5978$\pm$0.05 0.5428$\pm$0.03 0.6086$\pm$0.02 0.5700$\pm$0.01 0.6261$\pm$0.04 0.5016$\pm$0.02 0.5984$\pm$0.02
Ours vs. NLinUCB +0.3098 +0.3426 +0.2487 +0.2313 +0.2315 +0.1116 +0.1320 +0.1696 +0.2227 +0.1403 +0.2168
Table 2: Accuracy on the target domain for the VisDA17 dataset. Note that in this zero-shot target regret setting, the accuracy $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is the target regret.
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
Test Accuracy aeroplane bicycle bus car horse knife motorcycle person plant skateboard train truck Average
LinUCB [1] N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LinUCB-P [1] 0.0510$\pm$0.02 0.0190$\pm$0.00 0.1360$\pm$0.05 0.0440$\pm$0.02 0.0390$\pm$0.01 0.5580$\pm$0.06 0.0410$\pm$0.02 0.2050$\pm$0.05 0.0850$\pm$0.04 0.0920$\pm$0.13 0.0580$\pm$0.02 0.0460$\pm$0.12 0.1145$\pm$0.06
NLinUCB [11] 0.2870$\pm$0.03 0.0220$\pm$0.02 0.0000$\pm$0.00 0.1250$\pm$0.02 0.0000$\pm$0.00 0.0130$\pm$0.02 0.0470$\pm$0.02 0.0000$\pm$0.00 0.0460$\pm$0.02 0.0050$\pm$0.00 0.6550$\pm$0.04 0.0000$\pm$0.00 0.1001$\pm$0.02
NLinUCB-P [11] 0.0060$\pm$0.01 0.0020$\pm$0.01 0.9222$\pm$0.02 0.0010$\pm$0.00 0.0080$\pm$0.01 0.0990$\pm$0.03 0.0030$\pm$0.00 0.0020$\pm$0.00 0.0180$\pm$0.02 0.0420$\pm$0.01 0.0000$\pm$0.00 0.0150$\pm$0.01 0.0932$\pm$0.01
DABand (Ours) 0.5500$\pm$0.02 0.3560$\pm$0.04 0.4500$\pm$0.02 0.3850$\pm$0.01 0.5590$\pm$0.04 0.2650$\pm$0.03 0.6480$\pm$0.03 0.2410$\pm$0.03 0.4880$\pm$0.04 0.4060$\pm$0.02 0.7680$\pm$0.05 0.1970$\pm$0.02 0.4428$\pm$0.03
Ours vs. NLinUCB +0.2630 +0.3340 +0.4500 +0.2600 +0.5590 +0.2520 +0.6010 +0.2410 +0.4420 +0.4010 +0.1130 +0.1970 +0.3427
Table 3: Accuracy on the target domain for the S2RDA49 dataset. Note that in this zero-shot target regret setting, the accuracy $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is the target regret.
Metrics Acc Per Class $\uparrow$ Acc $\uparrow$
Test Accuracy aeroplane bicycle bus car knife motorcycle plant skateboard train truck Average
LinUCB [1] N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A
LinUCB-P [1] 0.1430$\pm$0.02 0.2600$\pm$0.03 0.0810$\pm$0.01 0.0240$\pm$0.01 0.1070$\pm$0.02 0.0570$\pm$0.01 0.1060$\pm$0.01 0.0390$\pm$0.01 0.0960$\pm$0.02 0.1260$\pm$0.03 0.1039$\pm$0.02
NLinUCB [11] 0.0350$\pm$0.00 0.1210$\pm$0.02 0.0510$\pm$0.01 0.0870$\pm$0.02 0.0270$\pm$0.00 0.0240$\pm$0.01 0.2720$\pm$0.04 0.0090$\pm$0.00 0.1730$\pm$0.02 0.0309$\pm$0.01 0.1108$\pm$0.02
NLinUCB-P [11] 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.2990$\pm$0.03 0.7000$\pm$0.02 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.0000$\pm$0.00 0.1160$\pm$0.01 0.1115$\pm$0.01
DABand (Ours) 0.5470$\pm$0.03 0.6410$\pm$0.03 0.5840$\pm$0.03 0.2140$\pm$0.01 0.0460$\pm$0.01 0.7570$\pm$0.04 0.5680$\pm$0.04 0.1340$\pm$0.02 0.1110$\pm$0.02 0.3120$\pm$0.02 0.3914$\pm$0.03
Ours vs. NLinUCB +0.5120 +0.5200 +0.5330 +0.1270 +0.0190 +0.7330 +0.2960 +0.1250 -0.0620 +0.2811 +0.2806
Predicted Reward. The predicted reward term in Eqn. 5, i.e., $\sum\nolimits_{i=1}^{N}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}$, can be translated to the loss term $\sum\nolimits_{i=1}^{N}\mathcal{L}_{i}^{reg}$, where for each round $i$ we have
$\mathcal{L}_{i}^{reg}=\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}\rangle\right|.$ (8)
Minimizing the predicted reward $\mathcal{L}_{i}^{reg}$ regularizes the model to avoid overestimating rewards.
Constant Term. In Eqn. 5, $\psi+C$ is the constant term, where $\psi$ is defined in Definition 3.7, and $C$ is a positive constant (see the Appendix for more details).
4.2 Putting Everything Together: DABand Training Algorithm
Putting all non-constant loss terms above together, we have the final minimax optimization problem
$\min\nolimits_{\widehat{\phi}}\max\nolimits_{g}\mathcal{L}^{DABand},$ (9)
where the value function (objective function):
${\mathcal{L}}^{DABand}=\sum\nolimits_{i=1}^{N}(\mathcal{L}_{i}^{\mathit{R}_{\mathcal{S}}}+\mathcal{L}_{i}^{\epsilon_{\mathcal{S}}(h)}+\mathcal{L}_{i}^{reg})+\lambda\cdot\mathcal{L}_{i}^{div}.$ (10)
Note that we need to alternate between updating $\widehat{\theta}$ using LinUCB [1] and updating the encoder $\widehat{\phi}$ with the discriminator $g$ in Eqn. 9.
Formally, our algorithm operates as follows: In each iteration, we access the context in the original feature space (i.e., $\mathbb{R}^{d}$) for each action from both source and target domains, denoted as $\left\{x_{i,a}^{\mathcal{S}}\right\}_{a\in[K]}$ and $\left\{x_{i,a}^{\mathcal{T}}\right\}_{a\in[K]}$ respectively. Subsequently, we compute the latent representation using the encoder $\phi(\cdot)$ and select the action $\widehat{a}_{i}$ for iteration $i$ based on the LinUCB selection rule. The bandit parameters (i.e., $\theta,A,b$) are updated in each iteration. The weights in $\phi$ are updated every episode of length $H$ (i.e., a batch with $H$ past iterations) using our objective function in Eqn. 9, following the same rule as in Neural-LinUCB [11]. Alg. 1 summarizes our algorithm.
5 Experiments
In this section, we compare DABand with existing methods on real-world datasets.
Datasets. To demonstrate the effectiveness of our DABand, We evaluate our methods in terms of prediction accuracy and zero-shot target regret on three datasets, i.e., DIGIT [37], VisDA17 [38], and S2RDA49 [39]. See details for each dataset in Appendix B.
Baselines. We compare our DABand with both classic and state-of-the-art contextual bandit algorithms, including LinUCB [1], LinUCB with principle component analysis (PCA) pre-processing (i.e., LinUCB-P), Neural-LinUCB [11] (NLinUCB), and Neural-LinUCB variant that incorporates PCA (i.e., NLinUCB-P). Details for each baselines and discussion are in Appendix C. Note that domain adaptation baselines are not applicable to our setting because it only works in offline settings and assumes complete observability of labels in the source domain (see Appendix C for details).
Zero-Shot Target Regret (Accuracy). We evaluate different methods on the zero-shot target regret setting in Definition 3.5, where all methods can only collect feedback from the source domain, but not the target domain. In this setting, accuracy is equal to $1$ minus the average target regret, i.e., $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is defined in Eqn. 2 and Definition 3.5. Therefore higher accuracy indicates lower target regret and better performance.
We report the per-class accuracy and average accuracy of different methods in Table 1, Table 2 and Table 3 for DIGIT, VisDA17 and S2RDA49, respectively. Our DABand demonstrates favorable performance against NLinUCB, a leading contextual bandit algorithm with representation learning. Despite NLinUCB’s superior representation learning capabilities through neural networks (NN) over LinUCB – yielding impressive performance in single-domain applications – its efficacy in cross-domain tasks is heavily contingent on the disparity between the source and target domains. Typically, the strengths of NN, which include enhanced representation learning (feature extraction), may inadvertently amplify this domain divergence, potentially leading to overfitting on the source domain. Conversely, our DABand algorithm not only improves accuracy but also adeptly addresses the challenges posed by domain divergence. Note that LinUCB encounters an out-of-memory issue due to the high-dimensionality of the data in VisDA17 and S2RDA49, marked as “N/A” in Table 2 and Table 3, respectively.
Figure 1: The cumulative regrets of different methods for 1,920 rounds on DIGIT (left), VisDA17 (middle), and S2RDA49 (right). Results are averaged over 5 runs. LinUCB is not reported for VisDA17 and S2RDA49 due to out-of-memory issues.
Continued Training in Target Domains and Cumulative Regret. Besides zero-shot target regret, we also evaluate how different methods perform when one continues their training in the high-cost target domain and starts to collect feedback (reward) after the model is trained in the source domain. A good cross-domain bandit model can provide a head-start in this setting, therefore enjoying significantly lower cumulative regret in the target domain and saving substantial cost in collecting feedback in the high-cost target domain.
Fig. 1 shows the cumulative regrets of different methods on DIGIT, VisDA17 and S2RDA49, respectively. We show results averaged over five runs with different random seeds. Our DABand consistently surpasses all baselines for all rounds in terms of both cumulative regrets and their increase rates, demonstrating DABand’s potential to significantly reduce the cost in high-cost target domains.
Table 4: Results of the ablation study in terms of accuracy (higher is better). Note that the accuracy $ACC=1-\tfrac{1}{N}R_{{\mathcal{T}}}$, where $R_{{\mathcal{T}}}$ is the target regret. “R” and “P” is short for “Regression Error” and “Predicted Reward”, respectively.
Datasets w/o R &\&& P w/o R w/o P DABand (Full)
DIGIT 0.5676 0.5716 0.5768 0.5984
VisDA17 0.4088 0.4238 0.4304 0.4642
S2RDA49 0.3691 0.3704 0.3719 0.3914
Ablation Study. To verify the effectiveness of each term in our DABand’s regret bound (Theorem 3.1, we report the accuracy of our proposed DABand after removing the Regression Error term and/or the Predicted Reward term (during training) in Table 4 for DIGIT, VisDA17 and S2RDA49. Results on all datasets show that removing either term will lead to performance drop. For example, in VisDA17, our full DABand achieves accuracy of $0.4642$; the accuracy drops to $0.4238$ after removing the Regression Error (R) term, and further drops to $0.4088$ after removing the Predicted Reward (P) term. These results verify the effectiveness of these two terms in our DABand.
6 Conclusion
In this paper, we introduce a novel domain adaptive neural contextual bandit algorithm, DABand, which adeptly combines effective exploration with representation alignment, utilizing unlabeled data from both source and target domains. Our theoretical analysis demonstrates that DABand can attain a sub-linear regret bound within the target domain. This marks the first regret analysis for domain adaptation in contextual bandit problems incorporating deep representation, shallow exploration, and adversarial alignment. We show that all these elements are instrumental in the domain adaptive bandit setting on real-world datasets. Moving forward, interesting future research could be to uncover more innovative techniques for aligning the source and target domains (mentioned in Appendix E), particularly within the constraints of: (1) bandit settings, as opposed to classification settings, and (2) the high-dimensional and dense nature of the target domain, contrasted with the sparse and simplistic nature of the source domain.
Appendix A Proofs
A.1 Proofs for Lemmas
Lemma A.1.
For any hypotheses $h,h^{\prime}\in\mathcal{H}$ and $a\in\mathcal{A}$ is selected by either $h$ or $h^{\prime}$,
$|\epsilon_{\mathcal{S}}(h,h^{\prime})-\epsilon_{\mathcal{T}}(h,h^{\prime})|\leq\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T}).$
Proof.
By definition of $\mathcal{H}\Delta\mathcal{H}$, we have
$\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$ $=2\enspace\sup_{h,h^{\prime}\in\mathcal{H}}|\mathbb{E}_{{\bf x}\sim\mathcal{S}}[|h({\bf x})-h^{\prime}({\bf x})|]-\mathbb{E}_{{\bf x}\sim\mathcal{T}}[|h({\bf x})-h^{\prime}({\bf x})|]|$
$=\dfrac{2}{N}\enspace\sup_{h,h^{\prime}\in\mathcal{H}}\Big{|}N\cdot\mathbb{E}_{{\bf x}\sim\mathcal{S}}[|h({\bf x})-h^{\prime}({\bf x})|]-N\cdot\mathbb{E}_{{\bf x}\sim\mathcal{T}}[|h({\bf x})-h^{\prime}({\bf x})|]\Big{|}$
$=\dfrac{2}{N}\enspace\sup_{h,h^{\prime}\in\mathcal{H}}\Big{|}\epsilon_{\mathcal{S}}(h,h^{\prime})-\epsilon_{\mathcal{T}}(h,h^{\prime})\Big{|}$
$\geq\dfrac{2}{N}\enspace\Big{|}\epsilon_{\mathcal{S}}(h,h^{\prime})-\epsilon_{\mathcal{T}}(h,h^{\prime})\Big{|}.$
Lemma A.2.
Given three hypotheses $h_{1},h_{2},h_{3}\in\mathcal{H}$, we have the following triangle inequality:
$\epsilon_{\mathcal{D}}(h_{1},h_{2})\leq\epsilon_{\mathcal{D}}(h_{1},h_{3})+\epsilon_{\mathcal{D}}(h_{2},h_{3}).$ (11)
Proof.
The hypothesis difference
$\epsilon_{\mathcal{D}}(h_{1},h_{2})$ $=\sum\nolimits_{i=1}^{N}\bigg{|}h_{1}(x^{\mathcal{D}})-h_{2}(x^{\mathcal{D}})\bigg{|}$
$=\sum\nolimits_{i=1}^{N}\bigg{|}h_{1}(x^{\mathcal{D}})-h_{2}(x^{\mathcal{D}})+h_{3}(x^{\mathcal{D}})-h_{3}(x^{\mathcal{D}})\bigg{|}$
$\leq\sum\nolimits_{i=1}^{N}\bigg{|}h_{1}(x^{\mathcal{D}})-h_{3}(x^{\mathcal{D}})\bigg{|}+\sum\nolimits_{i=1}^{N}\bigg{|}h_{3}(x^{\mathcal{D}})-h_{2}(x^{\mathcal{D}})\bigg{|}$
$=\epsilon_{S}(h_{1},h_{3})+\epsilon_{S}(h_{2},h_{3}).$
Lemma A.3.
Let $\mathcal{H}$ be a hypothesis space. Then for every $h\in\mathcal{H}$:
$\epsilon_{\mathcal{T}}(h)\leq\epsilon_{\mathcal{S}}(h)+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+\psi,$ (12)
where $\psi$ is defined in Definition 3.7.
Proof.
By definition in Definition 3.3 in main paper, we start from $\epsilon_{\mathcal{T}}(h)=\epsilon_{\mathcal{T}}(h,f_{\mathcal{T}})$, then we have
$\epsilon_{\mathcal{T}}(h)$ $=\epsilon_{\mathcal{T}}(f_{\mathcal{T}},h)$
$\leq\epsilon_{\mathcal{T}}(f_{\mathcal{T}},h^{*})+\epsilon_{\mathcal{T}}(h,h^{*})$
$\leq\epsilon_{\mathcal{T}}(h^{*})+\epsilon_{\mathcal{S}}(h,h^{*})+|\epsilon_{\mathcal{T}}(h,h^{*})-\epsilon_{\mathcal{S}}(h,h^{*})|$
$\leq\epsilon_{\mathcal{T}}(h^{*})+\epsilon_{\mathcal{S}}(h,h^{*})+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$
$\leq\epsilon_{\mathcal{T}}(h^{*})+\epsilon_{\mathcal{S}}(h)+\epsilon_{\mathcal{S}}(h^{*})+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$
$=\epsilon_{\mathcal{S}}(h)+\psi+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T}).$
A.2 Proof for the Target Regret Bound
Theorem A.1 (Target Regret Bound).
Denoting the contexts from the source domain ${\mathcal{S}}$ as $\{{\bf x}_{t,a}^{\mathcal{S}}\}_{i\in[N],a\in[K]}$, the associated ground-truth action as $\{a_{i}^{*}\}_{i=1}^{N}$, and the contexts from the target domain ${\mathcal{T}}$ as $\{{\bf x}_{t,a}^{\mathcal{T}}\}_{i\in[N],a\in[K]}$, the upper bound for our target regret $R_{{\mathcal{T}}}$ is
$\mathit{R}_{\mathcal{T}}$ $\triangleq\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta_{\mathcal{T}}^{*},\phi_{\mathcal{T}}^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\Bigg{)}$
$\leq{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}}_{Predicted~{}Reward}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}},$
where $\mathit{R}_{\mathcal{S}}$, $\epsilon_{\mathcal{S}}(h)$, and $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$ are defined in Sec. 3. $\psi$ and $C$ are constants.
Proof.
[Please check the full proof in the next page.]
By the definition of the Target Regret Bound, we have
$\mathit{R}_{\mathcal{T}}$ $=\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$=\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\epsilon_{\mathcal{T}}(h)+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle+\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\epsilon_{\mathcal{T}}(h)+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$=\epsilon_{\mathcal{T}}(h)+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\quad+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}-{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\leq\epsilon_{\mathcal{T}}(h)+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\quad+\Bigg{|}{\color[rgb]{0.5,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,1,0}\pgfsys@color@cmyk@stroke{0.50}{0}{1}{0}\pgfsys@color@cmyk@fill{0.50}{0}{1}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}-{\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}\Bigg{|}$
$\leq{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{T}}(h)}+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{T}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\quad+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle+\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\leq{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{S}}(h)+\psi+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}+\dfrac{N}{2}\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}C_{1}}$ (13)
$\quad+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}+{\color[rgb]{1,0.52,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.52,1}\pgfsys@color@cmyk@stroke{0}{0.48}{0}{0}\pgfsys@color@cmyk@fill{0}{0.48}{0}{0}\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}}$
$\leq\epsilon_{\mathcal{S}}(h)+\psi+N\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+C_{1}+\mathit{R}_{S}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}$
$\leq\epsilon_{\mathcal{S}}(h)+\psi+N\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+C_{1}+\mathit{R}_{S}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}+C_{2}$ (14)
$\leq\epsilon_{\mathcal{S}}(h)+\psi+N\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})+\mathit{R}_{S}+\sum^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\bigg{]}\Bigg{)}+C$ (15)
$={\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}}_{Predicted~{}Reward}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}},$
where $C=C_{1}+C_{2}$. Furthermore, the inequality (13) is due to the fact that the range of reward lies in $(0,1)$, which implies $\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}})\rangle-\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}})\rangle\Big{|}\bigg{]}\Bigg{)}\leq\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}1\bigg{]}\Bigg{)}:=C_{1}$, and the inequality (14) is due to $\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}})\rangle\Big{|}\bigg{]}\Bigg{)}\leq\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\theta^{*},\phi^{*}({\bf x}_{i,a_{i}^{*}})\rangle\Big{|}\bigg{]}\Bigg{)}:=C_{2}$. Furthermore, the inequality (15) is due to $\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}})\rangle\Big{|}\bigg{]}\Bigg{)}\leq\sum\nolimits^{N}_{i=1}\Bigg{(}\bigg{[}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}})\rangle\Big{|}\bigg{]}\Bigg{)}$. ∎
Appendix B Datasets
To evaluate the effectiveness of our DABand, for each dataset, we treat the “easy” domain as the low-cost source domain and the more challenging one as the high-cost target domain. This allows us to demonstrate the efficacy of our method by adapting from a simpler to a more complex domain within a contextual bandit setting.
DIGIT. Our DIGIT dataset consists of MNIST and MNIST-M. MNIST is a gray-scale hand-written digit dataset, while MNIST-M [37] features color digits. In this paper, MNIST serves as our source domain, and MNIST-M as our target domain, offering a more challenging adaptation path. The MNIST digits are gray-scale and uniform in size, aspect ratio, and intensity range, in stark contrast to the colorful and varied digits of MNIST-M. Therefore, adapting from MNIST to MNIST-M presents a greater challenge than adapting from MNIST-M to MNIST, where the target domain is less complex.
VisDA17. The VisDA-2017 image classification challenge [38] addresses a domain adaptation problem across 12 classes, involving three distinct datasets. The training set consists of 3D rendering images, whereas the validation and test sets feature real images from the COCO [40] and YouTube Bounding Boxes [41] datasets, respectively. Ground truth labels were provided only for the training and validation sets. Scores for the test set were calculated by a server operated by the competition organizers. For our purposes, we use the training set as the source domain and the validation set as the target domain.
S2RDA49. The S2RDA49 (Synthetic-to-Real) is a new benchmark dataset [39] constructed in 2023. This dataset contains $49$ classes. The source domain (i.e., the synthetic domain) is synthesized by rendering 3D models from ShapeNet [42]. The used 3D models are in the same label space as the target/real domain, and each class has 12K rendered RGB images. The target domain (i.e., the real domain) of S2RDA49 contains $60535$ images from $49$ classes, collected from the ImageNet validation set [43], ObjectNet [44], VisDA2017 validation set [38], and the web. In this paper, we select the only $10$ class that matches the VisDA17 dataset.
Appendix C Baselines
We compare our DABand with both classic and state-of-the-art contextual bandit algorithms. We start with LinUCB [1], a typical baseline for linear contextual bandit problems. To enable comprehensive comparisons, we extend our evaluation to include LinUCB augmented with pre-processed features using principle component analysis (PCA), referred to as LinUCB-P. In LinUCB-P, we first apply PCA to fit on training data (without labels/feedback) in both source and target domain. Then, with transformed, lower-dimensional context data, LinUCB is performed to see whether this pre-alignment procedure can transfer the knowledge to the target domain. Another pivotal baseline in our study is Neural-LinUCB [11] (NLinUCB), which utilizes several layers of fully connected neural networks to dynamically process the original features at the beginning of every iteration. Similar to LinUCB-P, we introduce a NLinUCB variant that incorporates PCA, i.e., NLinUCB-P. Note that domain adaptation baselines are not applicable to our setting. Specifically, domain adaptation methods only work in offline settings, and assume complete observability of labels in the source domain. In contrast, contextual bandit is an online setting where the oracle is revealed only when correctly predicted. Therefore domain adaptation methods are not applicable to our online bandit settings.
Appendix D Implementation Details
In this section, we provide detailed insights into the implementation of our approach, applied to two distinct datasets: DIGIT and VisDA17. We use Pytorch to implement our method, and all experiments are run on servers with NVIDA A5000 GPUs.
DIGIT. Within the DIGIT dataset framework, we use the MNIST dataset as the source domain and MNIST-M as the target domain. To ensure compatibility between the datasets, we standardize the channel size of images in the source domain ($c_{\mathcal{S}}=1$) to align with that in the target domain ($c_{\mathcal{T}}=3$). Each image undergoes normalization and is resized to $28\times 28$ pixels with 3 channels to accommodate the format requirements of both domains. Then, an encoder is utilized to diminish the data’s dimensionality to a more manageable latent space. Following this reduction, the data is processed through two fully connected neural network layers, ending in the final latent space necessary for loss computation as delineated in main paper. For the optimal hyperparameters, we set the learning rate to $1\times e^{-5}$, with $\lambda$ is chosen from $\{1.0,5.0,10.0,15.0,20.0\}$ and kept the same for all experiments. Additionally, we set the exploration rate $\alpha$ to $0.05$.
VisDA17. We use the VisDA17 dataset’s training set as the source domain, with the validation set functioning as the target domain. We adhere to preprocessing steps established by [45] to ensure uniformity across domains: Each image is normalized and resized to $224\times 224$ pixels with 3 channels, matching the requisite specifications for both source and target domains. For hyperparameters, the learning rate of $1\times e^{-5}$ is applied, with $\lambda$ is chosen from $\{1.0,5.0,10.0,15.0,20.0\}$ and then kept the same for all experiments. We set the exploration rate $\alpha$ to $0.05$.
S2RDA49. We selects $10$ classes from the original $49$ classes since it matches the target domain samples in VisDA17 [38]. We adhere to preprocessing steps established by [45] to ensure uniformity across domains: Each image is normalized and resized to $224\times 224$ pixels with 3 channels. For hyperparameters, the learning rate of $1\times e^{-3}$ is applied, with $\lambda$ chosen from $\{1.0,5.0,10.0,15.0,20.0\}$ and then kept the same for all experiments. We set the exploration rate $\alpha$ to $0.01$.
Appendix E Limitation
This work contains several limitations. Specifically, we highlight below:
Intuitive Perspective. The bandit algorithm indeed enjoys interpretability and the accuracy of its closed-form updates, but this is true only because it typically employs a linear model.
Unfortunately, linear models do not work very well for real-world data, which is often high-dimensional. For example, the empirical results for LinUCB and LinUCB-P in Table 1 and Table 2 show that such linear contextual bandit algorithms significantly underperform state-of-the-art neural bandit algorithms, which use back-propagation (similar results are shown in [9, 11]). In summary deep learning models with back-propagation is necessary due to the following reasons:
  • High-Dimensional Data. To enhance performance in real-world high-dimensional data (e.g., images), the integration of deep learning (and back-propagation) is necessary, though this may sacrifice a portion of the algorithm’s explanatory power.
  • Covariate Shift and Aligning Source and Target Domains in the Latent Space. In our settings where there is covariance shift between the source and target domains, a deep (nonlinear) encoder is required to transform the original context into a latent space where source-domain encodings and target-domain encoders can align. This also necessitates deep learning models with back-propagation.
Indeed, there is a trade-off between interpretability and performance. This would certainly be an interesting future direction, but it is out of the scope of this paper.
Theoretical Perspective. Our Theorem 3.1 is sharp, as all the inequalities are based on lemmas in the paper and the Cauchy inequality. Identifying the criteria under which the target regret bound reaches equality as well as how one can achieve it in practice would be interesting future work.
Empirical Perspective. The performance of DABand largely relies on the alignment quality between the source and target domains. Therefore, for two domains that cannot be aligned (for example, most domain adaptation tasks are predefined, and the data for both domains is pre-processed and cleaned, not original real-world data), it is challenging to evaluate how DABand can still transfer knowledge across different domains. Furthermore, it is still unknown whether, if we increase the domain shift in a dataset from another domain, our DABand can still perform well. These issues are beyond the scope of this paper, but they represent interesting areas for future work.
Appendix F Discussions
F.1 Difference between Bandit and Classification Settings
In the classification setting, given a sample ${\bf x}$, a model predicts its label $\widehat{y}$. Since its ground-truth label is known, one can directly apply the cross-entropy loss to perform back-propagation and update our model. In contrast, in bandit settings, for each sample ${\bf x}$, our model predicts an action $\widehat{a}$ by optimizing the estimated rewards. One then submits this predicted action to the environment and receives feedback that only indicates whether the predicted action $\widehat{a}$ is the optimal action $a^{*}$ or not. If not, we are informed that our prediction is incorrect, yet we do not receive information on what the correct (optimal) action should be. Moreover, during training, we only train on 300 episodes for DIGIT and VisDA17 and 100 episodes for S2RDA49. Each episode contains 64 samples. This implies that compared with those DA methods which train on multiple epochs, in our settings, all of the models only see a sample once. Such complexity makes the bandit setting much more challenging.
F.2 Importance of Bandits and DABand
Bandit algorithms are designed to navigate environments where feedback is sparse, costly, and indirect. By efficiently learning from limited feedback – identifying not just when a prediction is wrong but adapting without explicit guidance on the right choice – bandit algorithms offer a strategic advantage in dynamically evolving settings. Our DABand exemplifies this by leveraging a low-cost source domain to improve performance in a high-cost target domain, thereby reducing cumulative regret (improving accuracy) while minimizing operational costs. DABand not only reduces the expense associated with acquiring and labeling vast datasets but also capitalizes on the intrinsic adaptability of bandit algorithms to learn and optimize in complex, uncertain environments.
F.3 Discussion on Zero-shot Target Regret Bound
F.3.1 Significance of Theorem 3.1
Note that Theorem 3.1 is nontrivial. While it does resemble the generalization bound in domain adaptation, there are key differences. As mentioned in Observation (3) in Sec. 3.6, our target regret bound includes two additional crucial terms not found in domain adaptation. Specifically:
  • Regression Error in the Source Domain. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$, which defines the difference between the true reward from selecting action $\widehat{a}_{i}$ and the estimated reward for this action.
  • Predicted Reward. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$, which serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Moreover, we would also like to highlight the key novelty/challenges in deriving Theorem 3.1:
  • Generalization of $\mathcal{H}\Delta\mathcal{H}$ Distance to Regression. In the original multi-domain generalization bound [1], the $\mathcal{H}\Delta\mathcal{H}$ divergence between source and target domains is derived for classification models. However, contextual bandit is essentially a reward regression problem, and therefore necessitates generalizing the $\mathcal{H}\Delta\mathcal{H}$ distance from the classification case to the regression case. This presents a significant technical challenge, and leads to a series of modification in the proof.
  • Decomposing the Target Regret into the Source Regret with Other Terms. The subsequent challenge involves decomposing the regret bound for the target domain (i.e., the target regret) into a source regret term and other terms to upper-bound the target regret. This is necessary because we do not have access to feedback/reward from the target domain (we only have access to feedback/reward in the source domain). This process requires a nuanced understanding of the interplay between source and target domain dynamics within our model’s framework.
F.3.2 Testing Phase rather than Traditional Bandit Learning Settings
Our regret bound in Theorem 3.1 is a zero-shot regret bound for the target domain (with empirical results in the Zero-Shot Target Regret paragraph of Sec. 5), which corresponds to a testing phase.
However, we would like to clarify that extending this bound to handle continued training in the target domain (corresponding to the Continued Training in Target Domains and Cumulative Regret paragraph of Sec. 5) is straightforward. At a high level, we can have
$R_{\mathcal{T}}^{total}=R_{\mathcal{T}}^{zero-shot}+R_{\mathcal{T}}^{continued},$
with the first term handled by our DABand’s Theorem 3.1 and the second term handled by typical single-domain contextual bandit.
F.3.3 Scale of Source Regret and Predicted Rewards
Similar to Neural-LinUCB, in our DABand, the true reward is restricted to the range of $[0,1]$ (as we mentioned in Sec. 3.1); therefore it will not make the source regret unbounded.
Furthermore, our DABand algorithm tries to minimize the source regret while aligning the source and target domains. The minimization of the source regret is theoretically guaranteed, as discussed in recent work such as Neural-LinUCB [11].
Appendix G Potential Impact of DABand
Our DABand has many potential real-world applications, especially when obtaining responses is costly. For instance, in testing new drugs, we can construct responses from mice for new drug A, and then use DABand to obtain the zero-shot hypothetical regret bound for responses in humans. In another scenario, we can collect data (responses) on humans for another published, similar drug B, and then transfer knowledge by aligning the divergence between drug A and drug B. If both regrets reveal acceptable performance, we might not need excessive costs for back-and-forth testing, which significantly speeds up the process and reduces costs.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: It can be found in the Sec. 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Appendix E.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All the theoretical results and assumptions can be found in Sec. 3 and Appendix A, and all the proofs are in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes] .
  19. Justification: It is simple to replicate.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No] .
  24. Justification: Data are all available online, and We will release the code once this paper is accepted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All the things are in Appendix B, Appendix C and Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Table 1, Table 2 and Table 3 for DIGIT, VisDA17 and S2RDA49, respectively.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No] .
  39. Justification: All of our experiments don’t require high performance computing resource, and a single GPU with 16 GB memory is enough. Therefore we do not include in our paper.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We satisfy all the requirements.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Appendix G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This question is not applicable to our paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We follow the rule to mention and cite others’ assets used in this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This question is not applicable to our paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This question is not applicable to our paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This question is not applicable to our paper.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.